Assume a corporation has earnings before depreciation and taxes of $120,000, depreciation of $48,000, and that it has a 35 percent tax bracket.

a. Compute its cash flow using the following format. (Input all answers as positive values.)

Earnings before depreciation and taxes

Depreciation

Earnings before taxes

Taxes

Earnings after taxes

Depreciation

Cash Flow

b. How much would cash flow be if there were only $12,000 in depreciation? All other factors are the same.

c. How much cash flow is lost due to the reduced depreciation from $48,000 to $12,000?  ​
Answer #1
Case a Case b
Earnings before depreciation and taxes 120000 120000
Depreciation 48000 12000
Earnings before taxes 72000 108000
Taxes (72000*35%)=25200 (108000*35%)=37800
Earnings after taxes 46800 70200
Depreciation 48000 12000
Cash flow 94800 82200

Hence cash flow lost=(94800-82200)=$12600.